### v3

# This is the title

Abstract: We derive an uncertainty principle for Lipschitz maps acting on subsets of Banach spaces. We show that this nonlinear uncertainty principle reduces to the Heisenberg-Robertson-Schrodinger uncertainty principle for linear operators acting on Hilbert spaces.
Keywords: Uncertainty Principle, Lipschitz map, Banach space.
Mathematics Subject Classification (2020): 26A16, 46B99.
1. Introduction
Let $\mathcal{H}$ be a complex Hilbert space and AA be a possibly unbounded self-adjoint linear operator defined on domain $\mathcal{D}(A)\subseteq\mathcal{H}$. For $h\in\mathcal{D}(A)$ with $\|h\|=1$, define the uncertainty of AA at the point $h$ as
$\Delta_{h}(A)\coloneqq\|Ah-\langle Ah,h\rangle h\|=\sqrt{\|Ah\|^{2}-\langle Ah,h\rangle^{2}}.$
In 1929, Robertson [6] derived the following mathematical form of the uncertainty principle (also known as uncertainty relation) of Heisenberg derived in 1927 [4] (English translation of 1927 original article by Heisenberg). Recall that for two linear operators $A:\mathcal{D}(A)\to\mathcal{H}$ and $B:\mathcal{D}(B)\to\mathcal{H}$, we define $A,B]\coloneqq AB-BA$ and $\{A,B\}\coloneqq AB+BA$.
Theorem $1.$
[6, 1, 4, 9, 2, 5] (Heisenberg-Robertson Uncertainty Principle) Let $A:\mathcal{D}(A)\to\mathcal{H}$ and $B:\mathcal{D}(B)\to\mathcal{H}$ be self-adjoint operators. Then for all $h\in\mathcal{D}(AB)\cap\mathcal{D}(BA)$ with $\|h\|=1$, we have
(1) $\frac{1}{2}\left(\Delta_{h}(A)^{2}+\Delta_{h}(B)^{2}\right)\geq\frac{1}{4}\left(\Delta_{h}(A)+\Delta_{h}(B)\right)^{2}\geq\Delta_{h}(A)\Delta_{h}(B)\geq\frac{1}{2}|\langle[A,B]h,h\rangle|.$
In 1930, Schrodinger improved Inequality (1) [7] (English translation of 1930 original article by Schrodinger).
Theorem 1.2.
[7] (Heisenberg-Robertson-Schrodinger Uncertainty Principle) Let $A:\mathcal{D}(A)\to\mathcal{H}$ and $B:\mathcal{D}(B)\to\mathcal{H}$ be self-adjoint operators. Then for all $h\in\mathcal{D}(AB)\cap\mathcal{D}(BA)$ with $\|h\|=1$, we have
$\Delta_{h}(A)\Delta_{h}(B)\geq|\langle Ah,Bh\rangle-\langle Ah,h\rangle\langle Bh,h\rangle|=\frac{\sqrt{|\langle[A,B]h,h\rangle|^{2}+|\langle\{A,B\}h,h\rangle-2\langle Ah,h\rangle\langle Bh,h\rangle|^{2}}}{2}.$
Theorem 1.2 promotes the following question.
Question 1.3.
What is the nonlinear (even Banach space) version of Theorem 1.2?
In this short note, we answer Question 1.3 by deriving an uncertainty principle for Lipschitz maps acting on subsets of Banach spaces. We note that there is a Banach space version of uncertainty principle by Goh and Goodman [3] which differs from the results in this paper. It is interesting to note that the uncertainty principle of Game theory, derived by Székely and Rizzo is nonlinear [8].
2. Nonlinear Heisenberg-Robertson-Schrodinger Uncertainty Principle
Let $\mathcal{X}$ be a Banach space and $\mathcal{M}\subseteq\mathcal{X}$ be a subset such that $0\in\mathcal{M}$. Recall that the collection of all Lipschitz functions $f:\mathcal{M}\to\mathbb{C}$ satisfying $f(0)=0$, denoted by $\mathcal{M}^{\#}$ is a Banach space [10] w.r.t. the Lipschitz norm
$\|f\|_{\text{Lip}_{0}(\mathcal{M},\mathbb{C})}\coloneqq\sup_{x,y\in\mathcal{M},x\neq y}\frac{|f(x)-f(y)|}{\|x-y\|}.$
Let $\mathcal{M}\subseteq\mathcal{X}$ be a subset such that $0\in\mathcal{M}$. Let $A:\mathcal{M}\to\mathcal{X}$ be a Lipschitz map such that $A(0)=0$. Given $x\in\mathcal{M}$ and $f\in\mathcal{X}^{\#}$ satisfying $f(x)=1$, we define two uncertainties of AA at $(x,f)\in\mathcal{M}\times\mathcal{X}^{\#}$ as
$\Delta(A,x,f)\coloneqq\|Ax-f(Ax)x\|,$
$\nabla(f,A,x)\coloneqq\|fA-f(Ax)f\|_{\text{Lip}_{0}(\mathcal{M},\mathbb{C})}.$
Theorem 2.1.
(Nonlinear Heisenberg-Robertson-Schrodinger Uncertainty Principle) Let $\mathcal{X}$ be a Banach space and $\mathcal{M},\mathcal{N}\subseteq\mathcal{X}$ be subsets such that $0\in\mathcal{M}\cap\mathcal{N}$. Let $A:\mathcal{M}\to\mathcal{X}$, $B:\mathcal{N}\to\mathcal{X}$ be Lipschitz maps such that $A(0)=B(0)=0$. Then for all $x\in\mathcal{M}\cap\mathcal{N}$ and $f\in\mathcal{X}^{\#}$ satisfying $f(x)=1$, we have
$\frac{1}{2}\left(\nabla(f,A,x)^{2}+\Delta(B,x,f)^{2}\right)$ $\geq\frac{1}{4}\left(\nabla(f,A,x)+\Delta(B,x,f)\right)^{2}$
$\geq\nabla(f,A,x)\Delta(B,x,f)$
$\geq|f(ABx)-f(Ax)f(Bx)|.$
Proof.
$\nabla(f,A,x)\Delta(B,x,f)$ $=\|fA-f(Ax)f\|_{\text{Lip}_{0}(\mathcal{M},\mathbb{C})}\|Bx-f(Bx)x\|$
$\geq|[fA-f(Ax)f][Bx-f(Bx)x]|$
$=|f(ABx)-f(Bx)f(Ax)-f(Ax)f(Bx)+f(Ax)f(Bx)f(x)|$
$=|f(ABx)-f(Bx)f(Ax)-f(Ax)f(Bx)+f(Ax)f(Bx)\cdot 1|$
$=|f(ABx)-f(Ax)f(Bx)|.$
Corollary $2.$
Let $\mathcal{X}$ be a Banach space and $\mathcal{M},\mathcal{N}\subseteq\mathcal{X}$ be subsets such that $0\in\mathcal{M}\cap\mathcal{N}$. Let $A:\mathcal{M}\to\mathcal{X}$, $B:\mathcal{N}\to\mathcal{X}$ be Lipschitz maps such that $A(0)=B(0)=0$. Then for all $x\in\mathcal{M}\cap\mathcal{N}$ and $f\in\mathcal{X}^{\#}$ satisfying $f(x)=1$, we have
$\nabla(f,A,x)\Delta(B,x,f)+\|fB\|_{\text{Lip}_{0}(\mathcal{N},\mathbb{C})}\Delta(A,x,f)\geq|f([A,B]x)|.$
Proof.
Note that
$\nabla(f,A,x)\Delta(B,x,f)$ $\geq|f(ABx)-f(Ax)f(Bx)|$
$=|f(ABx-BAx)+f(BAx)-f(Ax)f(Bx)|$
$=|f([A,B]x)+(fB)[Ax-f(Ax)x]|$
$\geq|f([A,B]x)|-|(fB)[Ax-f(Ax)x]|$
$\geq|f([A,B]x)|-\|fB\|_{\text{Lip}_{0}(\mathcal{N},\mathbb{C})}\|Ax-f(Ax)x\|$
$=|f([A,B]x)|-\|fB\|_{\text{Lip}_{0}(\mathcal{N},\mathbb{C})}\Delta(A,x,f).$
Corollary 2.3.
Let $\mathcal{X}$ be a Banach space and $\mathcal{M},\mathcal{N}\subseteq\mathcal{X}$ be subsets such that $0\in\mathcal{M}\cap\mathcal{N}$. Let $A:\mathcal{M}\to\mathcal{X}$, $B:\mathcal{N}\to\mathcal{X}$ be Lipschitz maps such that $A(0)=B(0)=0$. Then for all $x\in\mathcal{M}\cap\mathcal{N}$ and $f\in\mathcal{X}^{\#}$ satisfying $f(x)=1$, we have
$\nabla(f,A,x)\Delta(B,x,f)+\|fB\|_{\text{Lip}_{0}(\mathcal{N},\mathbb{C})}\Delta(A,x,-f)\geq|f(\{A,B\}x)|.$
Proof.
Note that
$\nabla(f,A,x)\Delta(B,x,f)$ $\geq|f(ABx)-f(Ax)f(Bx)|$
$=|f(ABx+BAx)-f(BAx)-f(Ax)f(Bx)|$
$=|f(\{A,B\}x)-(fB)[Ax+f(Ax)x]|$
$\geq|f(\{A,B\}x)|-|(fB)[Ax+f(Ax)x]|$
$\geq|f(\{A,B\}x)|-\|fB\|_{\text{Lip}_{0}(\mathcal{N},\mathbb{C})}\|Ax-(-f)(Ax)x\|$
$=|f(\{A,B\}x)|-\|fB\|_{\text{Lip}_{0}(\mathcal{N},\mathbb{C})}\Delta(A,x,-f).$
Corollary 2.4.
(Functional Heisenberg-Robertson-Schrodinger Uncertainty Principle) Let $\mathcal{X}$ be a Banach space with dual $\mathcal{X}^{*}$, $A:\mathcal{D}(A)\to\mathcal{X}$ and $B:\mathcal{D}(B)\to\mathcal{X}$ be linear operators. Then for all $x\in\mathcal{D}(AB)\cap\mathcal{D}(BA)$ and $f\in\mathcal{X}^{*}$ satisfying $f(x)=1$, we have
$\frac{1}{2}\left(\nabla(f,A,x)^{2}+\Delta(B,x,f)^{2}\right)$ $\geq\frac{1}{4}\left(\nabla(f,A,x)+\Delta(B,x,f)\right)^{2}$
$\geq\nabla(f,A,x)\Delta(B,x,f)$
$\geq|f(ABx)-f(Ax)f(Bx)|.$
Corollary 2.5.
Let $\mathcal{X}$ be a Banach space, $A:\mathcal{D}(A)\to\mathcal{X}$ and $B:\mathcal{D}(B)\to\mathcal{X}$ be linear operators. Then for all $x\in\mathcal{D}(AB)\cap\mathcal{D}(BA)$ and $f\in\mathcal{X}^{*}$ satisfying $f(x)=1$, we have
$\nabla(f,A,x)\Delta(B,x,f)+\|fB\|\Delta(A,x,f)\geq|f([A,B]x)|.$
Corollary 2.6.
Let $\mathcal{X}$ be a Banach space, $A:\mathcal{D}(A)\to\mathcal{X}$ and $B:\mathcal{D}(B)\to\mathcal{X}$ be linear operators. Then for all $x\in\mathcal{D}(AB)\cap\mathcal{D}(BA)$ and $f\in\mathcal{X}^{*}$ satisfying $f(x)=1$, we have
$\nabla(f,A,x)\Delta(B,x,f)+\|fB\|\Delta(A,-x,f)$ $=\nabla(f,A,x)\Delta(B,x,f)+\|fB\|\Delta(A,x,-f)$
$\geq|f(\{A,B\}x)|.$
Corollary 2.7.
Theorem 1.2 follows from Theorem 2.1.
Proof.
Let $\mathcal{H}$ be a complex Hilbert space. $A:\mathcal{D}(A)\to\mathcal{H}$ and $B:\mathcal{D}(B)\to\mathcal{H}$ be self-adjoint operators. Let $h\in\mathcal{D}(AB)\cap\mathcal{D}(BA)$ with $\|h\|=1$. Define $\mathcal{X}\coloneqq\mathcal{H}$, $\mathcal{M}\coloneqq\mathcal{D}(A)$, $\mathcal{N}\coloneqq\mathcal{D}(B)$, $x\coloneqq h$ and
$f:\mathcal{H}\ni u\mapsto f(u)\coloneqq\langle u,h\rangle\in\mathbb{C}.$
Then
$\Delta(B,x,f)=\Delta(B,h,f)=\|Bh-f(Bh)h\|=\|Bh-\langle Bh,h\rangle h\|=\Delta_{h}(B),$
$\nabla(f,A,x)$ $=\nabla(f,A,h)=\|fA-f(Ah)f\|=\|fA-\langle Ah,h\rangle f\|$
$=\sup_{u\in\mathcal{H},\|u\|\leq 1}|f(Au)-\langle Ah,h\rangle f(u)|=\sup_{u\in\mathcal{H},\|u\|\leq 1}|\langle Au,h\rangle-\langle Ah,h\rangle\langle u,h\rangle|$
$=\sup_{u\in\mathcal{H},\|u\|\leq 1}|\langle u,Ah\rangle-\langle Ah,h\rangle\langle u,h\rangle|=\sup_{u\in\mathcal{H},\|u\|\leq 1}|\langle u,Ah-\langle Ah,h\rangle h\rangle|$
$=\|Ah-\langle Ah,h\rangle h\|=\Delta_{h}(A)$
and
$|f(ABx)-f(Ax)f(Bx)|$ $=|f(ABh)-f(Ah)f(Bh)|=|\langle ABh,h\rangle-\langle Ah,h\rangle\langle Bh,h\rangle|$
$=|\langle Bh,Ah\rangle-\langle Ah,h\rangle\langle Bh,h\rangle|=|\langle Ah,Bh\rangle-\langle Ah,h\rangle\langle Bh,h\rangle|.$
3. Acknowledgments
This research was partially supported by the University of Warsaw Thematic Research Programme “Quantum Symmetries”.